if a compressive force of 3*10^4 N is exerted on the end of 20cm long bone of cross sectional area 3.6cm², (a) will the bone break (b) if, not how much will it shorten? Given, compressive strength of bone =7.7*10^8 and Young`s modulus of bone =1.5*10^10N/m

Grade:11

1 Answers

Vikas TU
7 years ago
As the compressive force is less than the compressive strength of bone that is 7.7*10^8 N
Hence the bone would not break however it would have some strain.
Strain = stress/Y => F/AY => 3*10^4/(1.5*10^10*3.6 x 10^-4) = 0.005555 = delta(l)/l
delta(l) = 0.00555555*l => 0.005555*20/100 = 0.00111 meter. would be the final strain.